# Classification of three-qubit genuine entangled states using concurrence fill

### v1

@@ Abstract
Despite the successful experimental generation and verification of genuine multipartite entanglement, several existing entanglement measures remain insufficient to reliably capture its presence. In this study, we overcome this challenge by utilizing a geometric measure known as concurrence fill, which quantifies genuine multipartite entanglement using the area associated with the underlying concurrence triangle. Firstly, the concurrence fill for three-qubit pure states is reformulated using three tangle and partial tangles. This yields a representation that is more tractable and operational. Using this formulation, we derive a criterion to classify GHZ and W class of states. Next, we analyse the rank-2 mixture of generalized GHZ and W class of states. Furthermore, we derive the concurrence fill for the eigenstates of the corresponding mixture and obtain an upper bound for mixed states with zero tangle.
Keywords: Quantum Entanglement. Genuine entangled states. Concurrence fill
@@ I Introduction
Quantum entanglement is an important physical resource in quantum information theory, which distinguishes between quantum and classical worlds. Entanglement has applications in bipartite as well as multipartite quantum systems. Genuine multipartite entanglement (GME) is a significant concept for multipartite systems, since it plays an essential role in quantum communication karlson1998 ; yeo2006 and quantum cryptography hillery1999 . Recognizing the significance of entanglement as a valuable resource, it becomes natural to explore its quantification through theoretical methods. For the bipartite system, various entanglement measures have been developed, such as concurrence wootters1998 ; wootters1997 , negativity vidal2002 , entanglement of formation wootters1998 , etc. Carvalho et al. derived the generalization of concurrence, namely the generalized multipartite concurrence carvalho2004 . A novel entanglement monotone for genuine multipartite entanglement is proposed, serving as a natural extension of the negativity and coinciding with it in the bipartite case jungnitch2011 . Additionally, other bipartite entanglement measures, including the relative entropy of entanglement and the robustness of entanglement, are shown to be readily adaptable to the multipartite framework vidal2001 .
@@
Acin et al introduced the classification of mixed three-qubit states into different entanglement classes acin2001 . These states are classified into separable, biseparable, W class and GHZ class of states. Any nontrivial tripartite entangled state can be converted, by means of SLOCC, into one of two standard forms, namely, GHZ or W state, which are given as follows dur2000 :
$|GHZ\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$ (1)
$|W\rangle$ $=$ $\frac{1}{\sqrt{3}}\left(|100\rangle+|010\rangle+|001\rangle\right)$ (2)
@@
GHZ and W class of states are stochastically inequivalent classes of entangled states, i.e., a state in GHZ class cannot be converted into W class by stochastic LOCC.
@@
A tripartite measure of entanglement known as three-tangle (or tangle) was introduced by Coffman, Kundu and Wootters (CKW) as residual entanglement ckw2000 . Tangle was derived as consequence of the most significant property of the quantum entanglement, i.e., its monogamous behavior. It plays a significant role in the quantum cryptography which restricts the information gain by an eavesdropper qaisar2017 , and in quantum channel discrimination sen2016 . Monogamy relations were first established by Coffman, Kundu and Wootters (CKW) in ckw2000 by considering a three-qubit system. It was shown that the entanglement between the subsystems $A$ and $B$ and entanglement between the subsystems $A$ and $C$ cannot be greater than the entanglement between the subsystems $A$ and $BC$. The concept of monogamy property has been extended to the general $N$-qubit system verstrate2006 . Monogamy inequalities for various entanglement measures such as concurrence ckw2000 , entanglement of formation olive2014 , negativity fan2007 ; kim2009 , Tsallis-q entanglement kim2010 ; luo2016 and Renyi-$\alpha$ entanglement sanders2010 have been further studied.
@@
Quantification of entanglement for multipartite systems is a great challenge due to the much richer mathematical structures involved as compared to the bipartite case. A valid measure of multipartite entanglement must meet two key criteria to be considered a genuine multipartite entanglement (GME) measure. These criteria, identified by Ma et al. ma2011 , are: (a) the measure must be zero for all product and biseparable states, and (b) the measure must be positive for all nonbiseparable states (GHZ and W classes for the three-qubit case). In eberly2021 , authors used the monogamy inequality on entanglement distribution in three-qubit systems to introduce a new measure of genuine tripartite entanglement called concurrence fill. It is a GME measure since it is non-zero for GHZ as well as W class of states. The concurrence fill was studied in the experiments of three-flavor neutrino oscillations li2022 . The extension of concurrence fill to four partite systems, is presented in mishra , which represents concurrence fill as a combination of areas of cyclic quadrilateral and triangle structures, arising from two distinct types of bipartitions.
@@
In our work, we study the concurrence fill for pure and mixed three-qubit states. Firstly, we provide an analytic reformulation of concurrence fill in terms of three-tangle and partial tangle. This formulation is constructive because it has been shown in literature that tangle as well as partial tangles can be measured experimentally datta2018 . Since tangle is zero for W class of states, it cannot classify W class and GHZ class of states. To overcome this problem we use our formulation of concurrence fill to devise a criterion that classifies GHZ and W class of states in three-qubit systems. In eberly2021 , authors have introduced the concurrence fill for three-qubit pure states. A corresponding measure for mixed states can be obtained using convex roof extension. This method involves minimizing the concurrence fill of a mixed state $\rho$ over all possible pure state decompositions of $\rho$. Hence, convex roof extension is a powerful but computationally challenging tool. In this work, we consider the family of rank-2 mixture of generalized GHZ and W states in three qubit system. Since all the elements in a decomposition are superpositions of the eigenstates of the given density matrix, the minimum value of concurrence fill will depend on all the possible concurrence fill values for those superpositions.We derive the concurrence fill for the eigenstates of the density matrices belonging to this family of rank-2 mixed states. Next, we deduce a lower bound of the concurrence fill of these eigenstates. Moreover, we derive an upper bound for the concurrence fill of mixed states in this family of states.
@@ II Entanglement Measures
Several entanglement measures have been proposed, particularly for the bipartite states. Yet, the situation is much more complicated for multipartite systems. If an N-partite state is neither separable nor biseparable, it is genuine $N$-partite entangled. Any genuine multipartite entanglement measure (GME), denoted by $E(\rho)$ for the state $\rho$ satisfies the following necessary conditions vedral .
@@
    $E(\rho)$ is independent under local unitary transformations
@@ II.1 Concurrence
Concurrence is a well-defined measure of entanglement. Let $\rho_{AB}$ be a density matrix describing a two-qubit pure or mixed state. Then the spin flipped density matrix $\tilde{\rho}_{AB}$ is defined as
@@
$\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ (5)
@@
where $\lambda_{i}^{\prime}s$ are the real eigenvalues of the positive semi-definite matrix ${\rho}_{AB}\tilde{\rho}_{AB}$. If the two qubit state $\rho_{AB}$ is a pure state, concurrence is defined as
$C_{AB}=2\sqrt{det\rho_{A}}$ (7)
where $\rho_{A}$ is the reduced state (trace of $\rho_{AB}$ over the qubit $B$).
@@ II.3 Partial Tangle
To further characterize entanglement in three-qubit systems, a quantity known as partial tangle was introduced in slee2005 . It is defined as follows.
$\tau_{ij}=\sqrt{C^{2}_{i(jk)}-C^{2}_{ik}},\;i\neq j\neq k\;\text{and}\;i,j,k=1,2,3$ (10)
@@ II.4 Concurrence fill
Xie and Eberly introduced a triangle measure of entanglement for three-qubit systems, named as, concurrence fill eberly2021 . It has the advantage of being nonincreasing under local quantum operations and classical communications (LQCC). This measure is based on the entanglement polygon inequality which states that one entanglement cannot increase the sum of the other two qian2018 ; zhu2015 , i.e.,
@@
where a subscript $i$ refers to the system’s $i$th qubit. A concurrence triangle is the geometrical interpretation of entanglement polygon inequality eberly2021 .
[…]
@@
Concurrence fill, denoted by $C_{F}$, is a genuine tripartite entanglement measure that represents the area of the concurrence triangle. For a three qubit state, $\rho_{ABC}$, concurrence fill is given as
$C_{F}(\rho_{ABC})=\left[\frac{16}{3}Q(Q-C^{2}_{A|BC})(Q-C^{2}_{B|AC})\right.\\ \left.\times(Q-C^{2}_{C|AB})\right]^{1/4},\\ \text{where}\quad Q=\frac{1}{2}\left(C^{2}_{A|BC}+C^{2}_{B|AC}+C^{2}_{C|AB}\right)$ (12)
@@
$C_{F}(|GHZ\rangle)=1\;\text{and}\;C_{F}(|W\rangle)=\frac{8}{9}$ (13)
$C_{F}(|GHZ\rangle)=1$ and Also, $C_{F}$ is monotonic under LQCC.
Concurrence fill can be generalized for mixed states via convex roof extension. It is given as
$C_{F}(\rho_{ABC})=min_{\{p_{i},|\psi_{i}\rangle\}}\sum_{i}{p_{i}C_{F}(|\psi_{i}\rangle})$ (14)
@@ III Classification of GHZ and W-class of states
In this section, we delve into the geometric entanglement measure known as concurrence fill for three-qubit pure states. We begin by expressing concurrence fill in terms of the tangle and partial tangle, demonstrating that this formulation remains invariant under different basis choices. Subsequently, we introduce a method to classify three-qubit states into the GHZ and W classes using concurrence fill. To lay the groundwork, let us first examine the basis representation of three-qubit states.
It is known that any three-qubit pure state can be written in canonical form, where the positive Schmidt coefficients contain all the information about the non-local properties of quantum states acin2001 . Canonical forms of three-qubit pure states are classified based on minimal number of local basis product states, i.e., minimal number of non-local parameters, needed for the specification of a state. In three qubit system, there are three inequivalent classes of local basis product states given as follows.
$\{|000\rangle,|001\rangle,|010\rangle,|100\rangle,|111\rangle\}$ (15)
$\{|000\rangle,|001\rangle,|100\rangle,|110\rangle,|111\rangle\}$
$\{|000\rangle,|001\rangle,|011\rangle,|100\rangle,|111\rangle\}$ (16)
$\{|000\rangle,|010\rangle,|100\rangle,|101\rangle,|111\rangle\}$
$\{|000\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle\}$
$\{|000\rangle,|010\rangle,|011\rangle,|110\rangle,|111\rangle\}$ (17)
$\{|000\rangle,|001\rangle,|011\rangle,|101\rangle,|111\rangle\}$
where the three sets of states in (16) and (17) are related by permutation.
Acin et. al have shown that any three-qubit pure state $|\psi\rangle$ can be expressed as acinprl2001
$|\psi_{ABC}\rangle=\lambda_{0}|000\rangle+\lambda_{1}e^{i\theta}|100\rangle+\lambda_{2}|101\rangle$
$+\lambda_{3}|110\rangle+\lambda_{4}|111\rangle$ (18)
where $\lambda_{i}\geq 0$, $\sum_{i=0}^{4}\lambda_{i}=1$ and $\theta\in[0,\pi]$.
Note that the pure state $|\psi\rangle$ in (18), belongs to the class of symmetric states in (17).
Now we express the concurrence fill of a three-qubit state in terms of three-tangle and partial tangle. Let us first consider the state $|\psi_{ABC}\rangle$ given in (18).
The concurrences and tangle between any two qubits are given by
$\tau_{\psi}$ $=$ $4\lambda_{0}^{2}\lambda_{4}^{2}$ (19)
$C^{2}_{AC}$ $=$ $4\lambda_{0}^{2}\lambda_{2}^{2}$ (20)
$C^{2}_{AB}$ $=$ $4\lambda_{0}^{2}\lambda_{3}^{2}$ (21)
$C^{2}_{BC}$ $=$ $4(\lambda_{2}^{2}\lambda_{3}^{2}-\lambda_{1}^{2}\lambda_{4}^{2})^{2}$ (22)
@@
$\tau^{2}_{ij}=C^{2}_{ij}+\tau_{\psi}\;\;\;\forall\;\;i,j=1,2,3$ (23)
After evaluating the relevant expressions, the concurrence fill of the state $|\psi_{ABC}\rangle$ can be expressed as
$C_{F}(|\psi_{ABC}\rangle)=\left(\frac{16}{3}\prod_{i<j}(\tau_{ij}^{2}-\frac{\tau_{\psi}}{2})\sum_{i<j}(\tau_{ij}^{2}-\frac{3\tau_{\psi}}{2})\right)^{1/4}$ (24)
for $i,j=1,2,3$
It is interesting to note that the above expression of the concurrence fill is independent of the choice of basis, i.e., for any basis given in (15), (16) and (17), area under the concurrence triangle remains same.
@@
Theorem. The following inequality holds for W-class of states.
@@
Using (24) we have,
$\begin{split}C_{F}(|\psi_{ABC}\rangle)=\bigg{(}\frac{1}{3}(\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})(2\tau_{BC}^{2}-\tau_{\psi})\\ (2C^{2}_{BC}+C^{2}_{AC}+C^{2}_{AB}+\tau^{2}_{AB}+\tau^{2}_{AC}+\tau_{\psi})\bigg{)}^{1/4}\end{split}$ (26)
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq\bigg{(}\frac{1}{3}(\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})\\ (\tau_{BC}^{2}-\tau_{\psi})\sum_{i<j}(\tau_{ij}^{2}+C^{2}_{ij})\bigg{)}^{1/4}\end{split}$ (27)
Using AM-GM inequality on above equation, we have
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq((\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})\\ (\tau_{BC}^{2}-\tau_{\psi})\prod_{i<j}(\tau^{2}_{ij}+C_{ij}^{2})^{1/3})^{1/4}\end{split}$ (28)
i.e.,
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq(\tau_{AB}^{2}+C^{2}_{AB})^{1/3}(\tau_{AC}^{2}+C^{2}_{AC})^{1/3}\\ (\tau^{2}_{BC}+C_{BC}^{2})^{1/12}(\tau_{BC}^{2}-\tau_{\psi})^{1/4}\end{split}$ (29)
Proof follows from the inequality (29). ∎
@@
Consider the three-qubit state $|\phi\rangle=c|001\rangle+d|010\rangle+f|100\rangle+e|011\rangle$, where $|c|^{2}+|d|^{2}+|f|^{2}+|e|^{2}=1$
@@
    $|\phi_{1}\rangle=d|010\rangle+f|100\rangle+\frac{1}{\sqrt{5}}|011\rangle$. Note that $\tau(|\phi_{1}\rangle)=e^{2}f^{2}\neq 0$ which implies that $|\phi_{1}\rangle$ is GHZ state. Now using the criteria given in above result, we see that the inequality (30) is violated which implies that $|\phi_{1}\rangle$ belongs to GHZ class.
    Figure 2: The blue and yellow lines represent the lhs and rhs of the inequality (30) for the states $|\phi_{1}\rangle$.
@@
    For $e=0$, $c=\frac{1}{\sqrt{5}}$, $d,f\neq 0$, $|\phi_{2}\rangle=c|001\rangle+d|010\rangle+f|100\rangle$. In this case, we have
    $[C_{F}(|\phi_{2}\rangle)]^{4}\geq k\;\tau^{2}_{BC}$ (31)
    Thus, using (25), $|\phi_{2}\rangle$ belongs to W class. Also, since the three-tangle $\tau(|\phi_{2}\rangle)=e^{2}f^{2}=0$, and three partial tangles are non-zero, the state $|\phi_{2}\rangle$ belongs to W class of states.
    Figure 3: The blue and yellow lines represent the lhs and rhs of the inequality (31) for the states $|\phi_{2}\rangle$.

### v2

@@ Abstract
Despite the successful experimental generation and verification of genuine multipartite entanglement, several existing entanglement measures remain insufficient to reliably capture its presence. In this study, we overcome this challenge by utilizing a geometric measure known as concurrence fill, which quantifies genuine multipartite entanglement of pure states using the area associated with the underlying concurrence triangle. Firstly, the concurrence fill for three-qubit pure states is reformulated using three tangle and partial tangles. This yields a representation that is more tractable and operational. Using this formulation, we derive a criterion to classify GHZ and W class of states. Next, we analyse the rank-2 mixture of generalized GHZ and W class of states for which three-tangle is known to vanish over a zero polytope. Furthermore, we derive the concurrence fill for the eigenstates of the corresponding mixture and obtain its upper bound for mixed states with zero tangle.
@@ I Introduction
Quantum entanglement is an important physical resource in quantum information theory, which distinguishes between quantum and classical worlds. Entanglement has applications in bipartite as well as multipartite quantum systems. Genuine multipartite entanglement (GME) is a significant concept for multipartite systems, since it plays an essential role in quantum communication karlson1998 ; yeo2006 and quantum cryptography hillery1999 . Recognizing the significance of entanglement as a valuable resource, it becomes natural to explore its quantification through theoretical methods. For the bipartite system, various entanglement measures have been developed, such as concurrence wootters1998 ; wootters1997 , negativity vidal2002 , entanglement of formation wootters1998 , etc. Carvalho et al. derived the generalization of concurrence, namely the generalized multipartite concurrence carvalho2004 . A novel entanglement monotone for GME is proposed, serving as a natural extension of the negativity and coinciding with it in the bipartite case jungnitch2011 . Additionally, other bipartite entanglement measures, including the relative entropy of entanglement and the robustness of entanglement, are shown to be readily adaptable to the multipartite framework vidal2001 .
@@
Acin et al introduced the classification of mixed three-qubit states into different entanglement classes acin2001 ; acinprl2001 . These states are classified into separable, biseparable, and genuine entangled states. There are two different kinds of genuine entangled states, namely, GHZ and W class of states dur2000 . Any nontrivial tripartite entangled state can be converted, by means of stochastic local operations and classical communication (SLOCC), into one of two standard forms, namely, GHZ or W state, which are given as follows:
$|GHZ\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right),$ (1)
$|W\rangle$ $=$ $\frac{1}{\sqrt{3}}\left(|100\rangle+|010\rangle+|001\rangle\right).$ (2)
@@
GHZ and W class of states are stochastically inequivalent classes of entangled states, i.e., a state in GHZ class cannot be converted into W class by SLOCC dur2000 .
@@
A tripartite measure of entanglement known as three-tangle (or tangle) was introduced by Coffman, Kundu and Wootters (CKW) as residual entanglement ckw2000 . Tangle was derived as consequence of the most significant property of the quantum entanglement, i.e., its monogamous behavior. It plays a significant role in the quantum cryptography which restricts the information gain by an eavesdropper qaisar2017 , and in quantum channel discrimination sen2016 . Monogamy relations were first established by considering a three-qubit system ckw2000 . It was shown that the entanglement between the subsystems $A$ and $B$ and entanglement between the subsystems $A$ and $C$ cannot be greater than the entanglement between the subsystems $A$ and $BC$. The concept of monogamy property has been extended to the general $N$-qubit system verstrate2006 . Monogamy inequalities for various entanglement measures such as concurrence ckw2000 , entanglement of formation olive2014 , negativity fan2007 ; kim2009 , Tsallis-q entanglement kim2010 ; luo2016 and Renyi-$\alpha$ entanglement sanders2010 have been further studied.
@@
Quantification of entanglement for multipartite systems is a great challenge due to the much richer mathematical structures involved as compared to the bipartite case. A valid measure of multipartite entanglement must meet two key criteria to be considered a GME measure. These criteria, identified by Ma et al. ma2011 , are: (a) the measure must be zero for all product and biseparable states, and (b) the measure must be positive for all nonbiseparable states (GHZ and W classes for the three-qubit case). In eberly2021 , authors used the monogamy inequality on entanglement distribution in three-qubit systems to introduce a new measure of genuine tripartite entanglement called concurrence fill. It is a GME measure since it is non-zero for GHZ as well as W class of states. The concurrence fill was studied in the experiments of three-flavor neutrino oscillations li2022 . The extension of concurrence fill to four partite systems, is presented in mishra , which represents concurrence fill as a combination of areas of cyclic quadrilateral and triangle structures, arising from two distinct types of bipartitions.
Although concurrence fill reflects rich geometric aspects of entanglement and bipartite concurrences, and serves as a useful GME measure, it has its own limitations. Ge et al have shown with the help of an example that concurrence fill is not monotonic under local operations and classical communication (LOCC) Ge2023 . Interestingly, we have reformulated the concurrence fill in terms of tangle and partial tangles. Since, it has also been proved that the three-tangle is LOCC mononotic, it adds to the relavance of our work Ge2023 . Moreover, Xie and Eberly derived the concurrence fill for pure three-qubit states eberly2021 . However, a closed formula for the mixed states GME including that for the concurrence fill in the case of mixed states remains unknown. In njp2008 , the authors present an explicit expression for the three-tangle corresponding to a generalized mixture of GHZ and W classes of states. It has been shown that, for a certain range of the mixing parameter, the three-tangle of a mixed state vanishes. Consequently, a significant subset of states within this family cannot be characterized using the three-tangle, as it fails to detect their entanglement. This limitation adds to the motivation of our study: we aim to quantify the entanglement present in the mixed states $\rho(p)$ with vanishing three-tangle. To this end, we employ the concurrence fill as an alternative measure to estimate the entanglement in these states.
@@
In our work, we study the concurrence fill for pure and mixed three-qubit states. Firstly, we provide an analytic reformulation of concurrence fill in terms of three-tangle and partial tangle. This formulation is constructive because it has been shown in literature that tangle as well as partial tangles can be measured experimentally datta2018 . Since tangle is zero for W class of states, it cannot classify W class and GHZ class of states. To overcome this problem we use our formulation of concurrence fill to devise a criterion that classifies GHZ and W class of states in three-qubit systems. In eberly2021 , authors have introduced the concurrence fill for three-qubit pure states. A corresponding measure for mixed states can be obtained using convex roof extension. This method involves minimizing the concurrence fill of a mixed state $\rho$ over all possible pure state decompositions of $\rho$. Hence, convex roof extension is a powerful but computationally challenging tool. In this work, we consider the family of rank-2 mixture of generalized GHZ and W states in three qubit system. Since all the elements in a decomposition are superpositions of the eigenstates of the given density matrix, the minimum value of concurrence fill will depend on all the possible concurrence fill values for those superpositions. We derive the concurrence fill for the eigenstates of the density matrices belonging to this family of rank-2 mixed states. Next, we deduce a lower bound of the concurrence fill of these eigenstates. Moreover, we derive an upper bound for the concurrence fill of mixed states with vanishing tangle in this family of states.
@@ II Entanglement Measures
Several entanglement measures have been proposed, particularly for the bipartite states. Yet, the situation is much more complicated for multipartite systems. If an $N$-partite state is neither separable nor biseparable, it is genuine $N$-partite entangled. Any genuine multipartite entanglement measure, denoted by $E(\rho)$ for the state $\rho$ satisfies the following necessary conditions vedral .
@@
    $E(\rho)$ is independent under local unitary transformations, i.e.,
@@ II.1 Concurrence
Concurrence is a well-defined measure of entanglement wootters1998 . Let $\rho_{AB}$ be a density matrix describing a two-qubit state. Then the spin flipped density matrix $\tilde{\rho}_{AB}$ is defined as
@@
$\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ (5)
@@
where $\lambda_{i}^{\prime}s$ are the real eigenvalues of the positive semi-definite matrix ${\rho}_{AB}\tilde{\rho}_{AB}$.
If $\rho_{AB}$ is a pure state, its concurrence is reduced to a simple expression as follows ckw2000 :
$C_{AB}=2\sqrt{det(\rho_{A})}$ (7)
where $det(.)$ represents the matrix determinant; and $\rho_{A}$ is the reduced state (trace of $\rho_{AB}$ over the qubit $B$).
There are several alternative formulations of concurrence for bipartite quantum states hororev2009 ; aggarwal2021 . These approaches may lead to distinct generalizations for multipartite systems ligao2024 ; zhu2020 . These generalizations offer insight into different facets of entanglement, for instance, how different subsystems are entangled with one another, or how entanglement is distributed within a single subsystem cornelio2013 .
@@ II.3 Partial Tangle
To further characterize entanglement in three-qubit systems, a quantity known as partial tangle was introduced in slee2005 . For a three qubit system described by the state $|\psi_{ABC}\rangle$, partial tangles are defined as follows.
$\tau_{ij}=\sqrt{C^{2}_{i|jk}-C^{2}_{ik}}$ (10)
for $i\neq j\neq k$ and $i$, $j$, $k$ in $\{A,B,C\}$.
@@ II.4 Concurrence fill
Xie and Eberly introduced a triangle measure of entanglement for three-qubit systems, named as, concurrence fill eberly2021 . This measure is based on the entanglement polygon inequality which states that one entanglement cannot increase the sum of the other two qian2018 ; zhu2015 , i.e.,
@@
for any distinct $i$, $j$, $k$ in $\{A,B,C\}$. A concurrence triangle is the geometrical interpretation of entanglement polygon inequality eberly2021 (Refer to Fig-1).
$C^{2}_{C|AB}$$C^{2}_{A|BC}$$C^{2}_{B|AC}$AABBCC
@@
Concurrence fill, denoted by $C_{F}$, is a genuine tripartite entanglement measure that represents the area of the concurrence triangle. For a three qubit state, $|\psi_{ABC}\rangle$, concurrence fill is given as
$C_{F}(|\psi_{ABC}\rangle)=\left[\frac{16}{3}Q(Q-C^{2}_{A|BC})(Q-C^{2}_{B|AC})\right.\\ \left.\times(Q-C^{2}_{C|AB})\right]^{1/4},\\ \text{where}\quad Q=\frac{1}{2}\left(C^{2}_{A|BC}+C^{2}_{B|AC}+C^{2}_{C|AB}\right).$ (12)
@@
$C_{F}(|GHZ\rangle)=1\;\text{and}\;C_{F}(|W\rangle)=\frac{8}{9}.$ (13)
Concurrence fill can be generalized for mixed states via convex roof extension.
@@ III Classification of GHZ and W-class of states
In this section, we delve into the geometric entanglement measure known as concurrence fill for three-qubit pure states. We begin by expressing concurrence fill in terms of the tangle and partial tangle. Subsequently, we introduce a method to classify three-qubit states into the GHZ and W classes using concurrence fill.
Theorem 1. The concurrence fill for a three-qubit state $|\psi_{ABC}\rangle$ can be expressed as follows.
$C_{F}(|\psi_{ABC}\rangle)=\left(\frac{16}{3}\sum_{i\neq j}(\tau_{ij}^{2}-\frac{3\tau_{\psi}}{2})\prod_{i\neq j}(\tau_{ij}^{2}-\frac{\tau_{\psi}}{2})\right)^{1/4}$ (14)
for $i\in\{A,B\};j\in\{B,C\}$.
Proof.
@@
$\tau^{2}_{ij}=C^{2}_{ij}+\tau_{\psi}\;\;\;\forall\;\;i,j=A,B,C.$ (15)
Using (8) in the above equation, we have
$C^{2}_{i|jk}=\tau_{ij}^{2}+C^{2}_{ik}$ (16)
for distinct $i$, $j$, $k$ in $\{A,B,C\}$. Let us consider the semi-perimeter of the concurrence triangle defined in (12).
$Q$ $=$ $\frac{1}{2}\left(C^{2}_{A|BC}+C^{2}_{B|AC}+C^{2}_{C|AB}\right)$ (17)
$=$ $\frac{3}{2}\tau_{\psi}+\left(C^{2}_{AB}+C^{2}_{BC}+C^{2}_{AC}\right)$ (18)
$=$ $\left(\tau^{2}_{AB}+\tau^{2}_{BC}+\tau^{2}_{AC}\right)-\frac{3}{2}\tau_{\psi}$ (19)
where (18) and (19) follows from (15) and (16).
Using (16) and (19), we have
$Q-C^{2}_{i|jk}=\tau^{2}_{jk}-\frac{\tau_{\psi}}{2}$ (20)
for distinct $i$, $j$, $k$ in $\{A,B,C\}$. Using (19) and (20), the concurrence fill for any three-qubit pure state defined in (12) can be reformulated as
$C_{F}(|\psi_{ABC}\rangle)=\left[\frac{16}{3}\left((\tau^{2}_{AB}+\tau^{2}_{BC}+\tau^{2}_{AC})-\frac{3}{2}\tau_{\psi}\right)\right.\\ \left.\times(\tau^{2}_{AB}-\frac{\tau_{\psi}}{2})(\tau^{2}_{BC}-\frac{\tau_{\psi}}{2})(\tau^{2}_{AC}-\frac{\tau_{\psi}}{2})\right]^{1/4}$ (21)
which is equivalent to (14). ∎
@@
Theorem 2. The following inequality holds for W-class of states.
@@
Let $|\psi_{ABC}\rangle$ be a three-qubit state. Consider the concurrence fill for the state $|\psi_{ABC}\rangle$ defined using tangles and partial tangles in Theorem1Theorem-1. Let us first reformulate the following terms in (21). Using (15), we obtain
$\tau^{2}_{ij}-\frac{\tau_{\psi}}{2}=C^{2}_{ij}+\frac{\tau_{\psi}}{2}=\frac{\tau^{2}_{ij}+C^{2}_{ij}}{2}$ (23)
and applying (23), we get
$(\tau^{2}_{AB}+\tau^{2}_{BC}+\tau^{2}_{AC})-\frac{3}{2}\tau_{\psi}=\sum_{i\neq j}\frac{\tau^{2}_{ij}+C^{2}_{ij}}{2}$ (24)
for $i\in\{A,B\};j\in\{B,C\}$.
Using (23) and (24) in (21), we have,
$\begin{split}C_{F}(|\psi_{ABC}\rangle)=\bigg{(}\frac{1}{3}(\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})\\ (2\tau_{BC}^{2}-\tau_{\psi})\sum_{i\neq j}(\tau^{2}_{ij}+C^{2}_{ij})\bigg{)}^{1/4}.\end{split}$ (25)
Using AM-GM inequality on the three terms in $\sum_{i\neq j}(\tau_{ij}^{2}+C^{2}_{ij})$ in the above equation, we obtain
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq((\tau_{AB}^{2}+C^{2}_{AB})(\tau_{AC}^{2}+C^{2}_{AC})\\ (2\tau_{BC}^{2}-\tau_{\psi})\prod_{i\neq j}(\tau^{2}_{ij}+C_{ij}^{2})^{1/3})^{1/4}.\end{split}$ (26)
After re-arranging the terms, the inequality in (26) can be re-expressed as follows.
$\begin{split}C_{F}(|\psi_{ABC}\rangle)\geq(\tau_{AB}^{2}+C^{2}_{AB})^{1/3}(\tau_{AC}^{2}+C^{2}_{AC})^{1/3}\\ (\tau^{2}_{BC}+C_{BC}^{2})^{1/12}(2\tau_{BC}^{2}-\tau_{\psi})^{1/4}\end{split}$ (27)
Let $|\psi_{ABC}\rangle$ belongs to the W class of state, i.e., $\tau_{\psi}=0$, and hence the inequality (27) is reduced to (22) which proves the theorem. ∎
@@
Consider the three-qubit state $|\psi_{cdfe}\rangle=c|001\rangle+d|010\rangle+f|100\rangle+e|011\rangle$, where $|c|^{2}+|d|^{2}+|f|^{2}+|e|^{2}=1$
@@
    $|\psi_{1}\rangle=d|010\rangle+f|100\rangle+\frac{1}{2}|011\rangle$. Note that $\tau(|\psi_{1}\rangle)=e^{2}f^{2}\neq 0$ which implies that $|\psi_{1}\rangle$ is GHZ state. Now using the criteria given in Theorem2Theorem-2, we see that the inequality (28) is not violated (refer to Fig-2) which implies that $|\psi_{1}\rangle$ belongs to GHZ class.
    Figure 2: The blue and yellow lines represent the lhs and rhs of the inequality (28) for the states $|\psi_{1}\rangle$. The $x$-axis represents the state parameter $f$.
@@
    For $e=0$, $c=\frac{1}{\sqrt{5}}$, $d,f\neq 0$, $|\psi_{2}\rangle=c|001\rangle+d|010\rangle+f|100\rangle$. In this case, we have
    $C_{F}(|\psi_{2}\rangle)]^{4}\geq k\;\tau^{2}_{BC}.$ (29)
    Thus, using (22), $|\psi_{2}\rangle$ belongs to W class (Refer to Fig-3). Also, since the three-tangle $\tau(|\psi_{2}\rangle)=e^{2}f^{2}=0$, and three partial tangles are non-zero, the state $|\psi_{2}\rangle$ belongs to W class of states.
    Figure 3: The blue and yellow lines represent the lhs and rhs of the inequality (29) for the states $|\psi_{2}\rangle$. The $x$-axis represents the state parameter $f\in[0,\frac{2}{\sqrt{5}}]$.